# Astrometry meets Pulsar Timing Arrays: Synergies for Gravitational Wave Detection

### v1

@@ 3.1 Building the optimal estimator: the case of astrometry
$C^{\delta\vec{\theta}}_{ij}=R\cdot C_{ij}^{\delta\vec{x}}\cdot R^{T}\,.$ (3.5)
Keeping this in mind, we now derive general expressions for our estimator of the coefficients $p_{n}$ formally appearing in the sum (3.1), later specialising to the monopole and dipole case with $p_{0}$ and $p_{1}$. The determination of these coefficients allows us to infer the properties of the SGWB. [^7] In principle, these basis coefficients can also be spherical harmonics coefficients or individual pixels used to discretize the SGWB intensity map. Our analysis represents a special case where we only expand the SGWB intensity up to the dipole term, assuming a known dipole direction. A quadratic estimator for $p_{n}$ can be formally expressed as
$\hat{p}_{n}=\delta n^{i}_{a}E_{ab,n}^{ij}\delta n^{j}_{b}-b_{n}\,,$ (3.6)
in terms of a matrix $E_{ab,n}$ and a vector $b_{n}$. We wish to determine the corresponding values of these quantities which render the estimator unbiased, and minimizes its variance.
Following the procedure developed in [99] (see also [100]), we find
[…]
Thus, the quantity $b_{n}$ should be chosen as $b_{n}={\rm Tr}[{\boldsymbol{E}}_{n}{\boldsymbol{N}}]$ to ensure that the estimator is unbiased. We now calculate the variance of estimator (3.6), in order to minimise it. We impose $W_{nn}=1$ as overall normalisation. The variance is given by
$\mathrm{Cov}(p_{n},p_{n^{\prime}})=2{\rm Tr}[{\boldsymbol{C}}{\boldsymbol{E}}_{n}{\boldsymbol{C}}{\boldsymbol{E}}_{n^{\prime}}]$ (3.8)
The matrix ${\boldsymbol{E}}_{n}$ is [^8] In the noise dominated limit with the diagonal noise, the solution reduces to ${\boldsymbol{E}}_{n}={\boldsymbol{H}}_{n}/{\rm Tr}[{\boldsymbol{H}}_{n}{\boldsymbol{H}}_{n}]$, and $W_{nn^{\prime}}={\rm Tr}[{\boldsymbol{H}}_{n}{\boldsymbol{H}}_{n^{\prime}}]/{\rm Tr}[{\boldsymbol{H}}_{n}{\boldsymbol{H}}_{n}]$
${\boldsymbol{E}}_{n}=\frac{{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}{\boldsymbol{C}}^{-1}}{{\rm Tr}[{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}]}\,.$ (3.9)
Hence the unbiased, optimal estimator for $p_{n}$ is obtained by inverting the resulting matrix $W_{nn^{\prime}}$:
$\widetilde{p}_{n}=\sum_{n^{\prime}}[W]^{-1}_{nn^{\prime}}\hat{p}_{n^{\prime}}\,,$ (3.10)
while the corresponding inverse covariance matrix is given by
$\mathrm{Cov}^{-1}(\widetilde{p}_{n},\widetilde{p}_{n^{\prime}})=\frac{1}{2}{\rm Tr}[{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n}{\boldsymbol{C}}^{-1}{\boldsymbol{H}}_{n^{\prime}}]\,.$ (3.11)
In fact, this quantity corresponds to the Fisher matrix for the parameters $p_{n}$, assuming they are Gaussian distributed. I.e. we consider a likelihood $L$ of the form
$-2\ln L=\delta n^{i}_{a}[C^{-1}]^{ij}_{ab}\delta n^{j}_{b}+\ln\det{\boldsymbol{C}}\,+n\ln 2\pi.$ (3.12)
Specialising to the case where $p_{0}$ and $p_{1}$ correspond to monopole and dipole as in eq (3.1), in the noise dominated limit with $N_{ab}^{ij}=\sigma^{2}\delta_{ab}\delta^{ij}$, we obtain the Fisher matrix $\mathcal{F}$ for the estimators
$\mathcal{F}=\frac{1}{2\sigma^{4}}\begin{pmatrix}{\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{0}]\quad&{\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{1}]\\ {\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{1}]\quad&{\rm Tr}[{\boldsymbol{H}}_{1}{\boldsymbol{H}}_{1}]\end{pmatrix}\,.$ (3.13)
@@
Hence, in the noise-dominated regime, the Fisher matrix depends only on the ORF to the monopole and dipole of the SGWB, discussed in Section 2. The error in determining the individual parameters $p_{i}$ we are interested in is given by $\Delta p_{i}=\sqrt{[\mathcal{F}^{-1}]_{ii}}$.
@@
$\mathcal{F}_{0}$ $\equiv{\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{0}]\simeq 3\times N_{\mathrm{star}}^{2}$ (3.14)
$\mathcal{F}_{1}$ $\equiv{\rm Tr}[{\boldsymbol{H}}_{1}{\boldsymbol{H}}_{1}]\simeq 0.65\times N_{\mathrm{star}}^{2}$ (3.15)
@@
In the same limit, the cross terms depending on ${\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{0}]$ vanish in eq (3.13). In practice, to perform Fisher forecasts by means of Eq (3.13) we focus on the frequency domain, performing a frequency binning with $\Delta f=1/T_{\rm obs},$ with $T_{\rm obs}=15\;\rm years$, and summing over the Fisher matrices at each frequency bin. We relate the value of the GW energy density $\Omega_{\rm GW}$ to the intensity through equation (2.6). We assume for the latter a power-law ansatz $I(f)\propto f^{3-\gamma}$, and report our results in Fig 6 for different values of spectral slope, also indicating the detection threshold indicated by the IPTA joint analysis [9].
@@
The noise in the frequency domain is given by $\sigma^{2}_{f}=2\sigma^{2}T_{\rm cad}$. We take different values of $\sigma$ as shown in Figure 6 and the observational cadence $T_{\rm cad}=\mathrm{year}/15$, corresponding approximately to the observational cadence of Gaia [36].
@@
Figure 6: Forecasts for the magnitude of the SGWB energy density and dipolar anisotropy, as measured with astrometry. See the discussion after eq (3.15). Left: The error $\Delta\Omega_{\rm GW,0}$ associated with the monopole estimator $p_{0}$. Right: $\Delta\Omega_{\beta}$ associated with the dipole estimator $p_{1}$. We define $\Omega_{\beta}=\beta(4-n_{\Omega})\Omega_{\rm GW,0}$.
@@
Under the assumptions leading to eqs (3.14) and (3.15) we find that astrometric surveys with $0.01$ milli-arcsecond (mas) precision which monitor $N>10^{6}$ stars may be competitive with PTA experiments, in terms of sensitivity to the SGWB. The detection of the kinematic dipole, whose amplitude is suppressed by a factor $\beta$ relative to the monopole, will accordingly require about $10^{3}$ times more stars. We should note, though, that for the green line shown in fig. 6, the weak signal approximation likely breaks down, as we go towards a higher and higher number of stars ($N_{\rm star}\gtrsim 10^{6}$). The noise level $\sigma=0.01$ mas roughly corresponds to the expected astrometric accuracy that will be achieved by Gaia DR5 for the brightest objects in the survey. [^9] https://www.cosmos.esa.int/web/gaia/science-performance Thus, the forecast sensitivity with this value of the noise is very unlikely to be achieved in real Gaia DR5 data but may be possible with future missions like Theia [48, 50, 49].
@@ 3.2 Forecasts: astrometry in synergy with PTA
$B_{pa}$ $=$ $\frac{K_{p,a}I_{f}}{4\pi f}\,,$ (3.19)
$D_{ab}$ $=$ $H_{ab}I_{f}+\delta_{ab}\sigma^{2}_{a}\,.$ (3.20)
The tensor $\gamma_{pq}$ corresponds the standard Hellings-Downs inter-pulsar correlation between pulsar time delays, while $K_{p,a}$ and $H_{ab}$ denote the pulsar-star and star-star correlations studied in section 2. In general, the covariance matrices depend explicitly on the SGWB intensity $I_{f}$ evaluated at the frequency bin under examination. (We use the same notation as [101, 102].)
@@ 3.2.1 Some concrete examples
To demonstrate in concrete how the cross-correlation of astrometric and PTA data can be useful to characterize the SGWB signal – even when the astrometric data alone are not very constraining – we start with a simple example. Case 1: We parameterize the SGWB intensity as
$I(f)=\frac{A_{\rm GW}^{2}}{2}\,\left(\frac{f}{f_{\rm ref}}\right)^{3-\gamma}\,,$ (3.24)
with $f_{\rm ref}=1/\rm year$. We choose PTA parameters so as to recover the level of constraints set by the IPTA joint analysis [9] with 15 years of observation time, roughly corresponding to $\log_{10}A_{\rm GW}=-14.6\pm 0.16,\,\gamma=13/3\pm 0.45$ at $95\%$ C.L. For the astrometric analysis, we take 1000 stars with identical noise level $\sigma^{2}_{S}=2[\Delta\theta_{\rm rms}]^{2}T_{\rm cad}$ with $\Delta\theta_{\rm rms}=0.001\,\rm mas$ and $T_{\rm cad}=\mathrm{year}/24$, i.e. with each star observed twice per month.
Although typical astrometric datasets have a larger number of stars, and at the same time larger noise levels, this example suffices to make our point. Both the pulsar and the star configuration is supposed to be distributed uniformly across the sky.
The signal strength, frequency binning (with bin width $\Delta f=1/T_{\rm obs}$) and noise levels are plotted in the left panel of fig. 7. In the right panel of the same figure we plot the Fisher forecast for the parameters $\log A_{\rm GW}$ and $\gamma$ within this setup. We learn that the joint PTA+Astrometric dataset does provide an improvement over the PTA-only constraints, even when the astrometric data alone is not too informative. Nearly identical results are obtained on changing the number of stars while at the same time rescaling the noise using
$\Delta\theta_{\rm rms}(N_{\rm star})=0.001\times\sqrt{N_{\rm star}/1000}\,.$ (3.25)
This relation is a consequence of the weak signal limit which holds approximately in this particular case, with the approximation getting better and better on increasing the overall noise level. We verified that eq. (3.25) holds for a reasonable number of stars – but can not do so for a number of stars beyond $N_{\rm star}\gtrsim 5000$, due to the increasing computational costs involved. (But see below for a method to handle such situations.) With $N_{\rm star}=10^{5}$, relation (3.25) gives a noise level of $\Delta\theta_{\rm rms}=0.01$ milli-arcseconds (mas).
Figure 7: Noise vs signal (left) + Fisher forecast (right) for the Case 1 example of section 3.2.1. Limits are $95\%$ C.L and correspond to the joint PTA+Astrometric forecast.
We proceed presenting two more examples, plotted in Figure 8, denoted with Case 2 and 3. They represent a case where the overall noise level for astrometry $(\sigma^{2})$ is reduced by a factor of 2 (Figure 8 left panel), and a situation keeping the same noise properties for the astrometry set, but with the PTA system in the strong-signal (noiseless) regime (Figure 8 right panel). The astrometric data alone do not set meaningful constraints on the SGWB properties: however, they are still able to improve the constraining power of the PTA data exploiting cross-correlations among data sets. The improvement can be significant if the PTA datasets are not in the strong signal regime; however, it is much smaller if that is indeed the case.
Figure 8: Fisher forecast for the lower astrometric noise level with same PTA noise as above (left) and noiseless PTA (right). These are Case 2 and 3 examples of section 3.2.1. Limits are $95\%$ C.L and correspond to the joint PTA+Astrometric forecast.

### v2

@@ 3.1 Astrometry only
$C^{\delta\vec{\theta}}_{ij}=R\cdot C_{ij}^{\delta\vec{x}}\cdot R^{T}\,$, (3.7)
which is the actual covariance matrix that will be used for calculations to follow. For zero mean, Gaussian distributed astrometric deflections, the likelihood $\mathcal{L}$ takes the form
$-2\ln\mathcal{L}=\delta n^{i}_{a}[C^{-1}]^{ij}_{ab}\delta n^{j}_{b}+\ln\det{\boldsymbol{C}}\,+n\ln 2\pi.$ (3.8)
How well we can determine the parameters that describe our signal is controlled by the Fisher matrix, defined as [99]
$\mathcal{F}_{ij}=\left\langle-\frac{\partial\ln\mathcal{L}}{\partial\theta_{i}\partial\theta_{j}}\right\rangle\,$, (3.9)
given in terms of the expectation value of second derivatives of $\ln{\cal L}$. The Fisher matrix sets a lower bound on the precision with which these parameters can be measured, $\Delta\theta_{i,\rm min}=\sqrt{[\mathcal{F}^{-1}]_{ii}}$. For the likelihood in (3.8), the Fisher matrix reads [99]
$\mathcal{F}_{ij}$ $=\frac{1}{2}{\rm Tr}\left[C^{-1}C_{,i}C^{-1}C_{,j}\right]\,$ (3.10)
where $C_{,i}$ denotes the derivative of $C$ w.r.t the parameter $\theta_{i}$. Using the definition of the covariance (3.1), we see that
$C_{,i}=\sum_{n}\frac{\partial p_{n}}{\partial\theta_{i}}{\boldsymbol{H}}_{n}\,.$ (3.11)
Specialising to the case where our signal parameters $\theta_{i}$ correspond to the monopole and dipole amplitude $p_{0}$ and $p_{1}$ as in eq (3.1), the Fisher matrix reads
$\mathcal{F}_{ij}=\frac{1}{2}{\rm Tr}\left[C^{-1}H_{i}C^{-1}H_{j}\right]\,$, (3.12)
where $i,j$ run from 0 to 1. In the noise dominated limit with a diagonal noise covariance matrix $N_{ab}^{ij}=\sigma^{2}_{N}\delta_{ab}\delta^{ij}$, we obtain the Fisher matrix $\mathcal{F}$ for $p_{0},p_{1}$ as
$\mathcal{F}=\frac{1}{2\sigma^{4}_{N}}\begin{pmatrix}{\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{0}]\quad&{\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{1}]\\ {\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{1}]\quad&{\rm Tr}[{\boldsymbol{H}}_{1}{\boldsymbol{H}}_{1}]\end{pmatrix}\,$. (3.13)
@@
Hence, in the noise-dominated regime, the Fisher matrix depends only on the ORF of the monopole and dipole of the SGWB, discussed in Section 2.
@@
$\mathcal{F}_{0}$ $\equiv{\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{0}]\simeq 3\times N_{\mathrm{star}}^{2}\,$, (3.14)
$\mathcal{F}_{1}$ $\equiv{\rm Tr}[{\boldsymbol{H}}_{1}{\boldsymbol{H}}_{1}]\simeq 0.65\times N_{\mathrm{star}}^{2}\,.$ (3.15)
@@
In the same limit, the cross terms depending on ${\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{1}]$ vanish in eq (3.13). In practice, to perform Fisher forecasts by means of Eq (3.13) we focus on the frequency domain, performing a frequency binning with $\Delta f=1/T_{\rm obs},$ with $T_{\rm obs}=15\;\rm years$, and summing over the Fisher matrices at each frequency bin. We relate the value of the GW energy density $\Omega_{\rm GW}$ to the intensity through equation (2.7). We assume for the latter a power-law ansatz $I(f)=I_{0}(f/f_{*})^{2-\gamma}$, and report our results in Fig 6 for different values of spectral slope, also indicating the detection threshold indicated by the IPTA joint analysis [9].
@@
The noise in the frequency domain is given by $\sigma^{2}_{N}(f)=2\sigma^{2}T_{\rm cad}$, where we have assumed a white noise component only. We take different values of $\sigma$ as shown in Figure 6 and the observational cadence $T_{\rm cad}=\mathrm{year}/15$, corresponding approximately to the observational cadence of Gaia [36], i.e. $T_{\rm cad}$ represents the time interval (in seconds) between successive observations of a particular star.
Assuming noise dominated covariance, the uncertainties for the intensity monopole amplitude $I_{0}$ and the dipole amplitude $I_{\beta}\equiv(1-n_{I})\beta I_{0}$  [^11] Note the relation $n_{\Omega}=n_{I}+3$ arising from (2.7)., take the following form
$\mathcal{F}_{I_{0}I_{0}}$ $=\frac{T_{\rm obs}}{2\sigma^{4}_{N}}\frac{\,{\rm Tr}[{\boldsymbol{H}}_{0}{\boldsymbol{H}}_{0}]}{(4\pi)^{2}}\sum_{f}(f/f_{*})^{2n_{I}}\Delta f\,$, (3.16)
$\mathcal{F}_{I_{\beta}I_{\beta}}$ $=\frac{T_{\rm obs}}{2\sigma^{4}_{N}}\frac{\,{\rm Tr}[{\boldsymbol{H}}_{1}{\boldsymbol{H}}_{1}]}{(4\pi)^{2}}\sum_{f}(f/f_{*})^{2n_{I}}\Delta f\,$, (3.17)
with $n_{I}=2-\gamma$. To obtain a scaling relation, we replace the discrete sum by an integral,
$\sum_{f}(f/f_{*})^{2n_{I}}\Delta f\approx\int_{f=1/T_{\rm obs}}^{1/T_{\rm cad}}(f/f_{*})^{2n_{I}}df\,.$ (3.18)
Given that $n_{I}=-7/3<0$  [^12] The fiducial value $n_{I}=-7/3$ adopted here corresponds to the theoretical expectation for supermassive black hole binaries in circular orbits [100]. , the integral is dominated by the lower frequencies and can be approximated (for $n_{I}\neq-1/2)$ by
$\left[\int_{f=1/T_{\rm obs}}^{1/T_{\rm cad}}(f/f_{*})^{2n_{I}}df\right]\approx\left[-\frac{f_{\rm min}}{2n_{I}+1}\left(\frac{f_{\rm min}}{f_{*}}\right)^{2n_{I}}\right]\,.$ (3.19)
Since, for a given parameter $\theta$, $\Delta\theta=\sqrt{[\mathcal{F}^{-1}]_{\theta\theta}}$, we obtain
$\Delta I_{0}$ $=\sqrt{\frac{2}{3|2n_{I}+1|}}\frac{4\pi\sigma^{2}_{N}}{N_{\rm star}}\left(\frac{T_{\rm obs}}{1\,\mathrm{year}}\right)^{n_{I}}\,$, (3.20)
$\Delta I_{\beta}$ $=\sqrt{\frac{2}{0.65|2n_{I}+1|}}\frac{4\pi\sigma^{2}_{N}}{N_{\rm star}}\left(\frac{T_{\rm obs}}{1\,\mathrm{year}}\right)^{n_{I}}\,$, (3.21)
where we used the fact that $f_{\rm min}=1/T_{\rm obs}$ and $f_{*}=1/\mathrm{year}$. Using the relation between the intensity and energy density (2.7), allows us to convert these expressions to uncertainties in $\Omega_{\rm GW}$. We find
$\Delta\Omega_{\rm GW,0(\beta)}$ $=1.6(3.4)\times 10^{-8}\sqrt{\frac{|2n_{I}+1|}{4/3}}\left(\frac{\sigma}{0.01}\right)^{2}\left(\frac{T_{\rm cad}}{\mathrm{year}/15}\right)\left(\frac{10^{6}}{N_{\rm star}}\right)\left(\frac{T_{\rm obs}}{15\,\mathrm{year}}\right)^{n_{I}}\,$, (3.22)
We emphasise that these relations hold only in the noise dominated (weak-signal) limit and for sufficiently red tilted SGWB and assuming that this tilt is known and not itself being inferred from the data.
@@
Figure 6: Forecasts for the magnitude of the SGWB energy density and dipolar anisotropy, as measured with astrometry. See the discussion after eq (3.15). Left: The error $\Delta\Omega_{\rm GW,0}$ associated with the monopole estimator $p_{0}$. Right: $\Omega_{\mathrm{GW},\beta}$ associated with the dipole estimator $p_{1}$. We define $\Omega_{\mathrm{GW},\beta}=\beta(4-n_{\Omega})\Omega_{\rm GW,0}$. The dashed lines correspond to $\Delta\Omega_{\rm GW,0}\approx 8.05\times 10^{-9}$, which is the value of the energy density with the NG15 results  [101], and $\Delta\Omega_{\mathrm{GW},\beta}\approx 3.3\times 10^{-11}$ .
@@
Under the assumptions leading to eqs (3.14) and (3.15) we find that astrometric surveys with $0.01$ milli-arcsecond (mas) precision which monitor $N>10^{6}$ stars may be competitive with PTA experiments, in terms of sensitivity to the SGWB. The detection of the kinematic dipole, whose amplitude is suppressed by a factor $\beta$ relative to the monopole, will accordingly require about $10^{3}$ times more stars. We should note, though, that for the green and blue lines shown in fig. 6, the weak signal approximation likely breaks down, as we go towards a higher and higher number of stars ($N_{\rm star}\gg 10^{6}$). The noise level $\sigma=0.01$ mas roughly corresponds to the expected astrometric accuracy that will be achieved by Gaia DR5 for the brightest objects in the survey. [^13] https://www.cosmos.esa.int/web/gaia/science-performance Thus, the forecast sensitivity with this value of the noise is very unlikely to be achieved in real Gaia DR5 data but may be possible with future missions like Theia [48, 50, 49].
@@ 3.2 Forecasts: astrometry in synergy with PTA
$B_{pa}$ $=$ $\frac{K_{p,a}I_{f}}{{(4\pi)^{2}}f}\,$, (3.26)
$D_{ab}$ $=$ $\frac{H_{ab}}{{4\pi}}I_{f}+\delta_{ab}\sigma^{2}_{a}\,.$ (3.27)
The tensor $\gamma_{pq}$ corresponds the Hellings-Downs (HD) inter-pulsar correlation between pulsar time delays [^15] For the HD correlations, we adopt the normalization convention of [84] where $\gamma_{pp}=4/3$., while $K_{p,a}$ and $H_{ab}$ denote the pulsar-star and star-star correlations studied in section 2. In general, the covariance matrices depend explicitly on the SGWB intensity $I_{f}$ evaluated at the frequency bin under examination. (We use the same notation as [102, 103].)
@@ 3.2.1 Some concrete examples
To demonstrate in concrete how the cross-correlation of astrometric and PTA data can be useful to characterize the SGWB signal – even when the astrometric data alone are not very constraining – we start with a simple toy example.
Case 1: We parameterize the SGWB intensity as
$I(f)=\frac{A_{\rm GW}^{2}}{2{f_{\rm ref}}}\,\left(\frac{f}{f_{\rm ref}}\right)^{2-\gamma}\,$, (3.31)
using the quantities $A_{\rm GW}$ and $\gamma$ to make comparison with PTA analyses simpler, [^16] We follow the conventions outlined in [102] to relate these quantities, where $h_{c}(f)\equiv\sqrt{2fI(f)}=A_{\rm GW}(f/f_{\rm ref})^{\alpha}$, with $\gamma\equiv 3-2\alpha$. with $f_{\rm ref}=1/\rm year$. We choose PTA parameters so as to recover the level of constraints set by the IPTA joint analysis [9] with 15 years of observation time, roughly corresponding to $\log_{10}A_{\rm GW}=-14.6\pm 0.16,\,\gamma=13/3\pm 0.45$ at $95\%$ C.L. For the astrometric analysis, we take 1000 stars with identical noise level $\sigma^{2}_{S}=2[\Delta\theta_{\rm rms}]^{2}T_{\rm cad}$ with $\Delta\theta_{\rm rms}=0.0002\,\rm mas$ and $T_{\rm cad}=\mathrm{year}/52$, i.e. with each star observed once per week.
Although these numbers are not representative of typical astrometric datasets, which have a larger number of stars, and at the same time larger noise levels, this toy example suffices to make our point and is computationally feasible without making any approximations for the Fisher matrix. Both the pulsar and the star configuration is supposed to be distributed uniformly across the sky.
Figure 7: Fisher forecast for the Case 1 (left) and Case 2 (right) examples of section 3.2.1.
In the left panel of fig. 7, we plot the Fisher forecast for the parameters $\log A_{\rm GW}$ and $\gamma$ within this setup. We learn that the joint PTA+Astrometric dataset does provide an improvement over the PTA-only constraints, even when the astrometric data alone is not too informative.
Case 2: We proceed by presenting one more example, plotted in the right panel of fig. 7, keeping the same noise properties for the astrometry set, but with the PTA system in the strong-signal (noiseless) regime. We see that the improvement in constraints with the addition of astrometry can be sizeable if the PTA datasets are not in the strong signal regime; however, it is much smaller if that is indeed the case.